7. What is the off speed for a plane with an acceleration of 4.2 m/s2 that can take off after traveling 300 yds?
Answer #1

Acceleration, a = m/s^2

Displacement, s = 300 yds = 274.32 m

Using Equations of Motion,

v = \sqrt{2gs} = \sqrt{2*4.2 *274.32 } = 48 m/s \rightarrow Answer